### v1

@@ 5 Numerical reconstructions and discussions
The uniqueness proof directly lends itself to a slice-by-slice reconstruction algorithm, which can be viewed as a multi-dimensional analogue of the algebraic reconstruction algorithm due to Curtis and Morrow [20] for square lattices. The detailed procedure is listed in Algorithm 1. When recovering the conductivity on $E_{t}$, it employs a basis of the quotient space $\ker T^{(t)}/\ker T^{(t-1)}$. Like Curtis-Morrow algorithm [20], each iteration of the algorithm involves only simple algebraic operations. One essential step per loop is to solve for the potential with the Cauchy data specified on the boundary $J_{t-1}^{\mathrm{S}}$. Similar to the Cauchy problem for the Laplace equation, the problem is expected to be ill-conditioned.
@@
4:for t=d1t=d-1 to $d\lceil\frac{n+1}{2}\rceil$ do
5:  Compute the kernel $K_{t}\coloneqq\ker\Lambda_{\gamma}(\partial D\setminus J_{t}^{\mathcal{S}},J_{t}^{\mathcal{S}})$.
@@
11:  Determine $\gamma|_{E_{t}}$ by solving the system
@@
We present numerical results for cubic lattices to illustrate the feasibility of the algorithm. The relevant numerical results are shown in Figs. 6 and 7 for cubic lattices of different sizes $n$. Note that Algorithm 1 initiates the reconstruction process from a single corner of the cubic lattice. However, numerically the error grows as the iteration proceeds. The displayed reconstructions combine the eight solutions obtained by starting from each of the eight corners, and for each edge in the graph, we select the solution corresponding to the closest corner.
The results in Figs. 6 and 7 show that the algorithm indeed can accurately recover the conductivity from the DtN matrix $\Lambda_{\gamma}$ (when the lattice size $n$ is not large), which corroborates the proof strategy. For each fixed $n$, the pointwise error increases towards the center (i.e., the central part has the largest error), exhibiting clearly a depth-dependent resolution: the conductivity on the edges near the corners can be more accurately recovered to recover than that are far away from the corners. Moreover, as the lattice size $n$ increases from 8 to 13, the maximum error increases dramatically, from $10^{-9}$ (for $n=8$) to $10^{0}$ (for $n=13$), and the dominant error occurs in the central part of the cubic lattice. Nonetheless, the reconstruction accuracy near the corners is still quite satisfactory. Note that the error arises purely from the round-off error of the floating point operations in the algorithm. This issue is a clear manifestation of the severe ill-posed nature of the discrete Calderón problem: the problem is exponentially ill-posed with respect to the lattice size $n$, and the phenomenon is well known also in the 2D case [19, 2]. Therefore, in the presence of data noise, to obtain reasonable conductivity reconstructions, incorporating suitable regularization is indispensable.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
$n=8$ $n=9$ $n=10$
Figure 6: The exact conductivity (top) and the logarithmic error $\log_{10}|\mathbf{e}_{\gamma}|$ (bottom) of the recovered conductivity for the three-dimensional cubic lattice.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
$n=11$ $n=12$ $n=13$
Figure 7: The exact conductivity (top) and the logarithmic error $\log_{10}|\mathbf{e}_{\gamma}|$ (bottom) of the recovered conductivity for the three-dimensional cubic lattice.

### v2

@@ 5 Numerical reconstructions and discussions
The uniqueness proof lends itself to a slice-by-slice reconstruction algorithm, which can be viewed as a multi-dimensional analogue of the algebraic reconstruction algorithm due to Curtis and Morrow [20] for square lattices. The detailed reconstruction procedure is listed in Algorithm 1. When recovering the conductivity on the edge set $E_{t}$, it employs a basis $\{\varphi_{i}\}$ of the quotient space $\ker T^{(t)}/\ker T^{(t-1)}$. Like the Curtis-Morrow algorithm [20], each iteration of the algorithm involves only simple algebraic operations. Note that Algorithm 1 starts the reconstruction process from one corner of the hypercubic lattice. However, numerically the reconstruction error grows as the iteration proceeds. The final reconstructions should combine the different solutions obtained by starting from each of the corners, and for each edge in the graph, we select the solution corresponding to the closest corner.
We briefly discuss the computational complexity of Algorithm 1, with the notation $O_{d}(\cdot)$ indicating the involved constant depending on $d$. The main computational cost of the algorithm lies in the following three steps: line 3 (kernel evaluation), line 5 (potential recovery) and line 9 (conductivity update). First, computing the kernel $K_{t}$ at line 3 involves $O_{d}(n)$ evaluations, each requiring $O_{d}((n^{d-1})^{3})=O_{d}(n^{3d-3})$ operations (e.g., via singular value decomposition), leading to an overall complexity $O_{d}(n^{3d-2})$. Second, the potential recovery at line 5 involves solving for the potential with the Cauchy data (i.e., both Dirichlet and Neumann data) specified on the boundary $J_{t-1}^{\mathrm{S}}$. This step requires $O_{d}(n)$ solves of (sparse) linear systems of size $O_{d}(n^{d})\times O_{d}(n^{d})$, leading to the total complexity $O_{d}(n^{1+3d})$ (by the standard direct solver without exploiting the sparse structure). Alternatively, by exploiting an upper triangular and sparse structure, this step can be reduced to $O_{d}(n^{d})$ per solve, with a total complexity $O_{d}(n^{1+d})$, but this procedure is numerically less stable. Third, solving for $\bm{\gamma}$ in each layer at line 9 involves $O_{d}(n)$ solves, each with complexity $O_{d}((n^{d-1})^{3})=O_{d}(n^{3d-3})$, resulting in a total complexity $O_{d}(n^{3d-2})$. This step can be accelerated in practice by partitioning the large linear system into smaller ones, following the idea in the proof of Theorem 4.1. In the numerical experiments below, the first step is more expensive than the other two steps. Moreover, the computation can be parallelized between different directions.
@@
4:for t=d1t=d-1 to $d\lceil\frac{n+3}{2}\rceil$ do
5:  Compute the kernel $K_{t}\coloneqq\ker\Lambda_{\gamma}(\partial D\setminus J_{t}^{\mathcal{S}},J_{t}^{\mathcal{S}})$, and construct a basis $\{\varphi_{i}\}_{i}$.
@@
11:  Determine $\gamma|_{E_{t}}$ by solving the linear system
@@
We present numerical results for cubic lattices to illustrate the feasibility of the algorithm. The relevant numerical results are shown in Figs. 6 and 7 for cubic lattices of different sizes $n$. The results show that the algorithm indeed can accurately recover the conductivity $\gamma$ from the DtN matrix $\Lambda_{\gamma}$ (when the lattice size $n$ is not large), which corroborates the proof strategy. For each fixed $n$, the pointwise error increases dramatically towards the center (i.e., the central part has the largest error), exhibiting clearly a depth-dependent resolution: the conductivity on the edges near the corners can be more accurately recovered than the ones that are far away from the corners. Moreover, as the lattice size $n$ increases from 8 to 12, the maximum error increases dramatically, from $10^{-9}$ (for $n=8$) to $10^{0}$ (for $n=12$), and the dominant error occurs in the central part of the cubic lattice. Nonetheless, the reconstruction accuracy near the corners is still quite satisfactory. Note that the error arises purely from the round-off error of the floating point operations when implementing the algorithm on computers. This issue is a clear manifestation of the severe ill-posed nature of the discrete Calderón problem: the problem is expected to be exponentially ill-posed with respect to the lattice size $n$, and the phenomenon is observed also in the 2D square lattice case [19, 2]. Therefore, in the presence of data noise, to obtain reasonable conductivity reconstructions, incorporating suitable regularization is indispensable.
These numerical results indicate that the discrete Calderón problem has at best logarithmic stability estimates. However, a rigorous analysis of the stability issue (with an explicit dependence on the grid size $n$ and dimensionality $d$) seems very challenging, which in the context of Algorithm 1 essentially boils down to precisely tracking the condition number of various submatrices (line 5 of the algorithm) arising from the slicing procedure. Numerically, the most unstable step lies at Step 5 of recovering the potential from the Cauchy data; Indeed, similar to the Cauchy problem for the Laplace equation, the problem is expected to be ill-conditioned. Fig. 8 presents the variation of the condition number $\kappa_{t}$ of the submatrices (defined as the ratio between the largest and the smallest singular values) arising in the potential determination step. The condition number $\kappa_{t}$ grows exponentially with $t$, which clearly suggests at best a logarithmic type stability estimate, but for each fixed $t$, $\kappa_{t}$ is observed to be nearly independent of $n$.
n=8n={8} n=10n={10} n=12n={12}
Figure 6: The exact conductivity ((top)) and the logarithm of the pointwise error $\log_{10}|\mathbf{e}_{\gamma}|$ ((bottom)) of the recovered conductivity for the three-dimensional cubic lattice.
n=14n={14} n=16n={16} n=18n={18}
Figure 7: The exact conductivity ((top)) and the logarithm of the pointwise error $\log_{10}|\mathbf{e}_{\gamma}|$ ((bottom)) of the recovered conductivity for the three-dimensional cubic lattice.
Refer to caption Refer to caption
Figure 8: The condition number $\kappa_{t}$ of the linear system involved in Step 5 of Algorithm 1, with different $t$ varying from $6$ to $33$. The results are for $d=3$.